# SOAC: Spatio-Temporal Overlap-Aware Multi-Sensor Calibration using Neural Radiance Fields

### 27 Nov 2023 v1

@@ 1 Introduction
Multi-sensor calibration plays a key role in autonomous systems as it ensures accuracy, reliability, and robustness in safety-critical tasks such as localization [6] and perception [22] in self-driving. In typical multi-sensor setups, the sensors are attached to a common rigid body where the spatial relationship between them can be obtained through a rigid transformation matrix. It is important to identify the exact values of those matrices to correctly exploit the data provided by the sensors. The process of finding these spatial transformations is called extrinsic calibration, which is a topic that has been and is still being heavily studied thanks to the increasing popularity of multi-sensor algorithms. In addition to extrinsic calibration, without an external synchronization system, it is also necessary to perform temporal calibration. With incorrect temporal calibration, performance on different tasks can be severely hindered. Even though some of the existing methods take temporal miscalibration into account [35, 29, 12], most of them assume that the sensors are correctly synchronized. Due to the importance of sensor calibration, a multitude of calibration solutions exist in the literature, as highlighted in the review from Li et al. [17] and as summarized in Table 1. These methods can be classified into two main categories: target-based methods and targetless methods.
@@
Figure 1: Method overview – SOAC is a novel multimodal spatio-temporal calibration method for cameras and LiDAR in the context of autonomous driving. By alternating the training of multiple implicit scenes (section 3.2) and sensors co-registration from these representations (section 3.3), SOAC achieves precise self-supervised calibration from raw data acquired in unconstrained urban environments.
@@
Targetless methods do not require manually placed targets and thus can be used on sequences captured without user intervention. This makes them more suitable for large-scale deployment. These approaches usually rely on shared information (i.e. overlap) between the different sensors, which can be of different modalities. Wang et al. [38] propose a correspondence between the reflectivity of the LiDAR scans and the grayscale intensity of the camera images. Other methods propose to find matches of specific features, like edges [45] or semantic classes [16].
@@
We take inspiration from the aforementioned works by exploiting the fully differentiable properties of the implicit scene representation to achieve spatial and temporal calibration. Different from existing methods [12, 48, 41], we propose to represent the scene by using multiple NeRFs akin to their corresponding sensor and advocate to alternate the optimization target between NeRF training and sensor calibration (i.e. Figure 1). Our method avoids overfitting the pose optimization to partial regions of the scene, resulting in a more robust and accurate calibration.
Table 1: Comparison of calibration methods.
Methods
@@
Feature-based Wang et al. [38] X X -
@@
SOAC (our method)
@@ 2 Related Work
With NeRF and the papers improving upon it [1, 26], the main focus was on the quality of novel view synthesis in addition to training and rendering speeds. However, since these approaches often validate their claims on carefully curated datasets, it is often assumed that the input poses corresponding to the data are already available and are accurate. However, in real-world situations, some or all the captured frames might be unposed or suffer from inaccuracies, hence, significantly impacting the quality of the final reconstruction result [19]. Therefore, many works later on attempted to tackle this issue through different formulations and adaptations of the overall optimization problem.
@@
To register an image with incorrect or no pose, iNeRF [44] proposes to use an already trained NeRF. It finds the pose that minimizes the photometric difference between the captured image and the rendered result from the model. By focusing on regions of interest, it is able to register unseen images with high precision. Using this idea as a basis, Loc-NeRF [21] combines Monte Carlo localization method [5] with the use of a pretrained NeRF as a map, to build a real-time global localization method. CROSSFIRE [25] takes advantage of the NeRF model’s flexibility to learn not only the radiance and density information of the map, but also a descriptor field. During the localization process, by iteratively matching the descriptors from the query image and the information given by the NeRF model, this method is able to provide high precision localization. Nevertheless, all these methods require training a NeRF from precise camera poses first before being able to localize new query images.
@@
The first method to leverage the fully differentiable nature of NeRF to optimize the input poses through backpropagation is NeRF– [40]. It proposes to optimize both the NeRF and the input poses by representing them as embeddings and show higher novel view synthesis quality when trained from noisy poses. BARF [19] improves upon this idea by adding a coarse-to-fine component to this method. It progressively liberates the frequencies of the input positional encoding to prevent the optimization from getting stuck in local minimum. SCNeRF [14] adds camera distortion estimation and uses a different 6-vector rotation formulation in the optimization, while SPARF [36] achieves pose optimization with sparse input views by relying on pixel matching and depth consistency.
While the aforementioned methods need an initial estimate of the camera poses, some recent methods completely remove the need for prior poses. NoPe-NeRF [2] uses an off-the-shelf monocular depth estimator (i.e. DPT [31]) to regularize relative poses between successive images. GNeRF [23] relies on adversarial learning to coarsely estimate the initial poses before refining them in a second phase. IR-NeRF [46] improves upon GNeRF by regularizing the implicit pose estimator with the unposed real images, increasing its robustness.
Although the NeRF-based pose optimization methods achieve reasonable scene reconstruction by recovering accurate camera poses, they are not suited for autonomous driving data as they are not handling multi-modal observations neither taking into account the rigidity constraint between multiple sensors mounted on a vehicle.
NeRF-based Calibration.
@@
Considering a multi-sensor rig, accurate calibration is crucial to exploit the data provided by the sensors together. NeRF-based calibration methods [12, 48, 41] take advantage of the rigid constraint between the sensors and the differentiable nature of NeRF to efficiently solve this challenging task. These methods have the advantage of being targetless and self-supervised, as they do not rely on an annotated training dataset. The idea is to use a NeRF as a common scene representation. Each sensor provides its observation (RGB images, depth measurement or point clouds), to both train the NeRF to represent the scene and to optimize its own extrinsic calibration parameters to fit the NeRF representation. In INF [48], the goal is to find the extrinsic transformation between a 360° camera to a LiDAR. First, the density network of NeRF is trained using the LiDAR depth data. Then, the whole scene’s radiance is trained using images, while simultaneously calibrating the camera. This method is limited to the calibration of a single 360° camera with a LiDAR, whereas autonomous driving systems rely on multiple cameras with narrower fields of view. AsyncNeRF [41] calibrates a pair of camera and depth sensors. It takes into account the temporal miscalibration between the sensors, by building a trajectory function. Nevertheless, the time offset is provided as input and not determined through optimization, which limits its utilization for spatio-temporal calibration. MOISST [12] proposes to accomplish temporal calibration in addition to extrinsic calibration, and to do so with any number of LiDARs and cameras, by training the NeRF with all the data, while also optimizing the prior extrinsic transformations and time offsets. By using a single NeRF to fuse the information from all the sensors, we cannot prevent degenerate cases where the estimation of the extrinsic parameters of one sensor diverges and causes the NeRF to learn a wrong scene geometry without correlating multi-sensor observations. Our method, SOAC, aims to achieve better robustness and calibration performance by leveraging the use of multiple NeRFs to counterbalance such limitations.
@@
(a) Scene representation training step – The parameters ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\hat{\Theta}}$ of each NeRF are trained with the images from their associated cameras and the LiDAR scans. The LiDAR calibration is also optimized through ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}T^{n_{(i+2)}}}$.
(b) Extrinsic and temporal optimization step – The real frame from the sensor is compared to the predicted frame on the other NeRFs to calculate the losses. The calibration is then optimized with backpropagation through the poses ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}T^{n_{(i+1)}}}$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}T^{n_{(i+2)}}}$.
Figure 2: Overview of the two training steps.
@@ 3 Method
Our multi-sensor calibration problem is formulated as follows: given a vehicle trajectory and initial priors of sensor positions mounted on the vehicle, we aim to recover the exact spatio-temporal calibration of the sensors on the vehicle. Our method is composed of two optimization steps that are performed sequentially all along the training (cf. Figure 1). The first step consists of training multiple implicit scene representations (NeRFs), one by camera, using only the observations from the dedicated sensor. During the second optimization step, we refine the extrinsic and temporal parameters of each sensor using the trained NeRF of all the other sensors in a round-robin manner. The motivation behind this design is to prevent over-fitting, calibration divergence, or implicit model convergence to a poor local minimum when all the observations are fused within the same implicit representation, as in MOISST [12].
@@ 3.1 Notations and Background
Our goal is to find the optimal transformations $\hat{\prescript{}{r}{T}^{i}}$ and time offsets $\hat{\delta_{i}}$ of the different sensors to calibrate with respect to the reference camera. The poses of the reference camera $r$ can be obtained by relying on an IMU, SLAM [27], or Structure-from-Motion [33]. Similar to MOISST, we build a continuous trajectory of the reference sensor $r$, $\mathcal{T}_{r}$, from the discrete poses of $r$ using linear interpolation for the pose translation and spherical linear interpolation (SLERP [34]) for the rotation. This trajectory is expressed as a function of time, that returns the pose of the reference camera $r$ for any given time $t$: $\prescript{}{w}{T}^{r}(t)=\mathcal{T}_{r}(t)$. Using the extrinsic transformations and the time offsets between the other sensors and the reference camera $r$, we can compute the absolute pose of sensor $i$ at specific timestamps with the following equation:
@@ 3.2 Scene Representation Training
For each camera sensor, a dedicated NeRF with parameters $\Theta_{i}$ is trained using rays that are generated exclusively from camera $i$. Each NeRF model with parameters $\Theta_{i}$ will only learn the part of the scene that is observed by its respective camera sensor $i$ (cf. Figure 1(a)). The color loss for training the scene representation is:
@@ 3.3 Extrinsic and Temporal Optimization
During the calibration step, our objective is to optimize the extrinsic transformation matrix $\prescript{}{r}{T}^{i}$ and temporal parameters $\delta_{i}$ by optimizing poses of camera $i$ using all the NeRFs, except the NeRF of parameters $\Theta_{i}$ associated to the current camera being calibrated (cf. Figure 1(b)). Using this optimization formulation, we enforce the images captured by each camera to be coherent with the NeRF trained by the other cameras. The camera calibration loss can be written as:
@@
and by considering equation 1, the optimization objective during the spatio-temporal optimization step is:
@@ 3.4 LiDAR Calibration
As a LiDAR only provides geometric information, we cannot register an RGB image to a NeRF which was only trained on LiDAR scans. This means that the registration step (cf. section 3.3) could not be accomplished on a LiDAR-trained NeRF. Instead of dedicating a NeRF for each LiDAR, we both train the camera NeRFs with all the LiDAR scans, and calibrate the LiDARs against all NeRFs (cf. Figure 2). Thus, we have for both the NeRF training step and calibration step:
@@
with ${D}^{n_{i}}$ the point cloud scan $n_{i}$ of LiDAR $i$. When adding the LiDAR loss in the objective equation 3, it becomes:
@@ 3.5 Visibility Grid
(a) Grid filling – Rays from camera $C_{i}$ fill the visibility grid linked to Nerf $\Theta_{i}$.
(b) Ray filtering – For cameras ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}C}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}j,\forall j}\neq{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}i}}$, rays are kept or filtered according to visibility from LABEL:sub@fig:filter_grid/filling
Figure 3: Illustration of visibility grid in practice.
In a multi-sensor setup, the portions of the scene observed from the different sensors might not overlap. This can lead to noisy reconstruction on the NeRF model if inference is performed at the unobserved regions (cf. Figure 4). To overcome this problem, NeRF2NeRF [10] performs pairwise registration of two NeRF models produced from different viewpoints by aligning the partially overlapping geometry of the two models. In a similar sense, we aim to consider the overlapping geometry from our different NeRFs that have been learned separately from each camera.
To achieve this, a boolean visibility grid for each NeRF model is reconstructed by considering the rays belonging to its akin sensor (see Figure 2(a)) during the scene representation step (Section 3.2). During the calibration step (Section 3.3) we exploit this visibility grid to only consider rays that overlap with trained regions on each NeRF used for registration (cf. Figure 2(b)). The grids are reinitialized every few epochs to account for the new poses resulting from the calibration refinements.
@@
Figure 4: Results using visibility grids on a Pandaset sequence – The predictions are done with the NeRF trained by the front camera.

### 27 Mar 2024 v3

@@ 1 Introduction
Multi-sensor calibration plays a key role in autonomous systems as it ensures accuracy, reliability, and robustness in safety-critical tasks such as localization [6] and perception [22] in self-driving. In typical multi-sensor setups, the sensors are attached to a common rigid body where the spatial relationship between them can be obtained through a rigid transformation matrix. It is therefore important to identify the exact values of those matrices to correctly exploit and merge the data provided by the sensors. The process of finding these spatial transformations is called extrinsic calibration, which is a topic that has been and is still being heavily studied thanks to the increasing popularity of multi-sensor algorithms. In addition to spatial calibration, without an external synchronization system, it is also necessary to perform temporal calibration. Using temporally miscalibrated sensors, performance on different tasks can be severely hindered. Although certain approaches in the literature address temporal misalignment [37, 30, 12], the prevailing assumption among these methods is the presence of properly synchronized sensors. Due to the importance of sensor calibration, a multitude of calibration solutions exist in the literature, as highlighted in the review from Li et al. [17] and summarized in Tab. 1. They can be classified into two main categories: target-based and targetless methods.
@@
Figure 1: Method overview. SOAC is a novel multimodal spatio-temporal calibration method for cameras and LiDAR in the context of autonomous driving. By alternating the training of multiple implicit scenes (Sec. 3.2) and sensors co-registration from these representations (Sec. 3.3), SOAC achieves precise self-supervised calibration from raw data acquired in unconstrained urban environments.
@@
Targetless methods do not require manually placed targets and thus can be used on sequences captured without user intervention. This makes them more suitable for large-scale deployment. These approaches usually rely on shared information (i.e. overlap) between the different sensors, which can be of different modalities. Wang et al. [40] and Pandey et al. [29] propose a correspondence between the reflectivity of the LiDAR scans and the grayscale intensity of the camera images. Other methods propose to find matches of specific features, like edges [47] or semantic classes [16].
@@
We take inspiration from the aforementioned works by exploiting the fully differentiable properties of the implicit scene representation to achieve spatial and temporal calibration. Different from existing methods [12, 50, 43], we propose to represent the scene by using multiple NeRFs akin to their corresponding sensor and advocate to alternate the optimization target between NeRF training and sensor calibration (i.e. Fig. 1). Our method avoids overfitting the pose optimization to partial regions of the scene, resulting in a more robust and accurate calibration.
@@
Feature-based Pandey et al. [29] X X -
@@
SOAC (ours)
Table 1: Comparison of calibration methods.
@@ 2 Related Work
With NeRF and the papers improving upon it [1, 26], the main focus was on the quality of novel view synthesis in addition to training and rendering speeds. However, since these approaches often validate their claims on carefully curated datasets, it is often assumed that the input poses corresponding to the data are already available and are accurate. However, in real-world situations, some or all the captured frames might be unposed or suffer from inaccuracies, hence, significantly impacting the quality of the final reconstruction result [19]. Therefore, several works later on attempted to tackle this issue through different formulations and adaptations of the overall optimization problem.
@@
To register an image with incorrect or no pose, iNeRF [46] proposes to use an already trained NeRF. It finds the pose that minimizes the photometric difference between the captured image and the rendered result from the model. By focusing on regions of interest, it is able to register unseen images with high precision. Using this idea as a basis, Loc-NeRF [21] combines Monte Carlo localization method [5] with the use of a pretrained NeRF as a map, to build a real-time global localization method. CROSSFIRE [25] takes advantage of the NeRF model’s flexibility to learn not only the radiance and density information of the map, but also a descriptor field. During the localization process, by iteratively matching the descriptors from the query image and the information given by the NeRF model, this method is able to provide high-precision localization. Nevertheless, all these methods require training a NeRF from precise camera poses first before being able to localize new query images.
@@
The first method to leverage the fully differentiable nature of NeRF to optimize the input poses through backpropagation is NeRF– [42]. It proposes to optimize both the NeRF and the input poses by representing them as embeddings and show higher novel view synthesis quality when trained from noisy poses. BARF [19] improves upon this idea by adding a coarse-to-fine component to this method. It progressively liberates the frequencies of the input positional encoding to prevent the optimization from getting stuck in a local minimum. SCNeRF [14] adds camera distortion estimation and uses a different 6-vector rotation formulation in the optimization, while SPARF [38] achieves pose optimization with sparse input views by relying on pixel matching and depth consistency. While the aforementioned methods need an initial estimate of the camera poses, some recent methods completely remove the need for prior poses. NoPe-NeRF [2] uses an off-the-shelf monocular depth estimator (i.e. DPT [32]) to regularize relative poses between successive images. GNeRF [23] relies on adversarial learning to coarsely estimate the initial poses before refining them in a second phase. IR-NeRF [48] improves upon GNeRF by regularizing the implicit pose estimator with the unposed real images, increasing its robustness. Although the NeRF-based pose optimization methods achieve reasonable scene reconstruction by recovering accurate camera poses, they are not suited for autonomous driving data as they do not handle multi-modal observations nor take into account the rigidity constraint between multiple sensors mounted on a vehicle.
NeRF-based Sensor Calibration.
@@
NeRF-based calibration methods [12, 50, 43] take advantage of the rigid constraint between the sensors and the differentiable nature of NeRF to efficiently solve this challenging task. These methods have the advantage of being targetless and self-supervised, as they do not rely on an annotated training dataset. The idea is to use the NeRF as a common scene representation. Each sensor provides its observations (RGB images, depth measurement, or point clouds), to both train the NeRF to represent the scene and to optimize its own extrinsic calibration parameters to fit the NeRF representation. In INF [50], the goal is to find the extrinsic transformation between a 360° camera and a LiDAR. First, the density network of NeRF is trained using the LiDAR depth data. Then, the whole scene’s radiance is trained using images, while simultaneously calibrating the camera. This method is limited to the calibration of a single 360° camera and a LiDAR, whereas autonomous driving systems rely on multiple cameras with narrower fields of view. AsyncNeRF [43] calibrates a pair of camera and depth sensors. It takes into account the temporal miscalibration between the sensors, by building a trajectory function. Nevertheless, the time offset is provided as input and not determined through optimization, which limits its utilization for spatio-temporal calibration. MOISST [12] proposes to accomplish temporal calibration in addition to extrinsic calibration, and to do so with any number of LiDARs and cameras, by training the NeRF with all the data, while also optimizing the prior extrinsic transformations and time offsets. By using a single NeRF to fuse the information from all the sensors, we cannot prevent degenerate cases where the estimation of the extrinsic parameters of one sensor diverges and causes the NeRF to learn a wrong scene geometry without correlating multi-sensor observations. Our method, SOAC, aims to achieve better robustness and calibration performance by leveraging the use of multiple NeRFs to counterbalance such limitations.
@@
Figure 2: SOAC training strategy. LABEL:sub@fig:method/step1 Scene representation training (Sec. 3.2): The parameters ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\hat{\Theta}}$ of each NeRF are trained with the images from their associated cameras and the LiDAR scans. The LiDAR calibration is also optimized through ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}T^{n_{(i+2)}}}$. LABEL:sub@fig:method/step2 Extrinsic and temporal optimization (Sec. 3.3): The real frame from the sensor is compared to the predicted frame on the other NeRFs to calculate the losses. The calibration is then optimized with backpropagation through the poses ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}T^{n_{(i+1)}}}$ and ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}T^{n_{(i+2)}}}$.
@@ 3 Method
Our multi-sensor calibration problem is formulated as follows: given a vehicle trajectory and initial priors of sensor poses mounted on the vehicle, we aim to recover the exact spatio-temporal calibration of the sensors on the vehicle. Our method is composed of two optimization steps that are performed sequentially all along the training (cf. Fig. 1). The first step consists of training multiple implicit scene representations (NeRFs), one by camera, using only the observations from the dedicated sensor. During the second optimization step, we refine the extrinsic and temporal parameters of each sensor using the trained NeRF of all the other sensors in a round-robin manner. The motivation behind this design is to prevent over-fitting, calibration divergence, or implicit model convergence to a poor local minimum when all the observations are fused within the same implicit representation, as in MOISST [12].
@@ 3.1 Notations and Background
Our goal is to find the optimal transformations $\hat{\prescript{}{r}{T}^{i}}$ and time offsets $\hat{\delta_{i}}$ of the different sensors with respect to the reference camera. The poses of the reference camera $r$ can be obtained by relying on IMU, SLAM [27], or Structure-from-Motion [35]. Similar to MOISST, we build a continuous trajectory of the reference sensor $r$, $\mathcal{T}_{r}$, from the discrete poses of $r$ using linear interpolation for the pose translation and spherical linear interpolation (SLERP [36]) for the rotation. This trajectory is expressed as a function of time, that returns the pose of the reference camera $r$ for any given time $t$: $\prescript{}{w}{T}^{r}(t)=\mathcal{T}_{r}(t)$. Using the extrinsic transformations and the time offsets between the other sensors and the reference camera $r$, we can compute the absolute pose of sensor $i$ at specific timestamps with the following equation:
@@ 3.2 Scene Representation Training
For each camera sensor, a dedicated NeRF with parameters $\Theta_{i}$ is trained using rays that are generated exclusively from camera $i$. Each NeRF model with parameters $\Theta_{i}$ will only learn the part of the scene that is observed by its respective camera sensor $i$ (cf. Fig. 1(a)). The color loss for training the scene representation is:
@@ 3.3 Extrinsic and Temporal Optimization
During the calibration step, our objective is to optimize the extrinsic transformation matrix $\prescript{}{r}{T}^{i}$ and temporal parameters $\delta_{i}$ by optimizing the poses of camera $i$ using all the NeRFs, except the NeRF of parameters $\Theta_{i}$ associated to the current camera being calibrated (cf. Fig. 1(b)). Using this optimization formulation, we enforce the images captured by each camera to be coherent with the NeRF trained by the other cameras. The camera calibration loss can be written as:
@@
and by considering Eq. 1, the optimization objective during the spatio-temporal optimization step is:
@@ 3.4 LiDAR Calibration
As LiDARs only provide geometric information, we cannot register an RGB image to a NeRF which was only trained on LiDAR scans. This means that the registration step (cf. Sec. 3.3) could not be accomplished on a LiDAR-trained NeRF. Instead of dedicating a NeRF for each LiDAR, we simultaneously train the camera NeRFs with all the LiDAR scans, and calibrate the LiDARs against all NeRFs (cf. Fig. 2). Thus, we have for both the NeRF training step and calibration step:
@@
with ${D}^{n_{i}}$ the point cloud scan $n_{i}$ of LiDAR $i$. When adding the LiDAR loss in the objective Eq. 3, it becomes:
@@ 3.5 Visibility Grid
Figure 3: SOAC’s visibility grid (Sec. 3.5). LABEL:sub@fig:filter_grid/filling Grid filling: Rays from camera $C_{i}$ fill the visibility grid linked to Nerf $\Theta_{i}$. LABEL:sub@fig:filter_grid/filtering Ray filtering: For cameras ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}C}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}j,\forall j}\neq{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}i}}$, rays are kept or filtered according to visibility from LABEL:sub@fig:filter_grid/filling.
In a multi-sensor setup, the NeRF representation exploits the overlap between sensors w.r.t. the whole sequence rather than a particular frame as for traditional targetless methods. However, the portions of the scene observed from the different sensors might not entirely overlap. This can lead to noisy reconstruction in the NeRF model if inference is performed at the unobserved regions (cf. Fig. 4). To overcome this problem, NeRF2NeRF [10] performs pairwise registration of two NeRF models produced from different viewpoints by aligning the partially overlapping geometry of the two models. In a similar sense, we aim to consider the overlapping geometry from our different NeRFs that have been learned separately from each camera.
To achieve this, a boolean visibility grid for each NeRF model is reconstructed by considering the rays belonging to its akin sensor (see Fig. 2(a)) during the scene representation step (Sec. 3.2). During the calibration step (Sec. 3.3) we exploit this visibility grid to only consider rays that overlap with trained regions on each NeRF used for registration (cf. Fig. 2(b)). The grids are reinitialized every few epochs to account for the new poses resulting from the calibration refinements.
@@
Figure 4: Visualization of the visibility grid (Sec. 3.5). Predictions done with the NeRF trained by front camera on a Pandaset [44] sequence.